OpenStudy (anonymous):

simultaneous equations a+3b=1 2a-b=1

OpenStudy (anonymous):

There are 2 simple ways to simple ways to solve these system of equations, 1: elimination, 2: substitution. Which method do you want to use?

OpenStudy (anonymous):

I would recommend using elimination. Start by multiplying the bottom equation by 3.

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

yeh soz

OpenStudy (anonymous):

a+3b=1 2a-b=1 Start by finding a or b. like: a=1-3b then input that value as a in the 2nd one. 2(1-3b)-b=1 Then calculate that until you find a valvue for b, then use b to find a value for a.

OpenStudy (anonymous):

got it thanks